# Implicit High-Order Moment Tensor Estimation and Learning Latent Variable Models

### v1

@@ 3.1 Recursive Pseudo-projections and their Properties
We start by defining the notion of a pseudo-projection.
@@
At any point in the circuit defining $\mathcal{S}$ if there is an order-$s$ tensor, we can compute the first $s$ steps of $\Phi$ applied to that value. The goal is to for every wire in the circuit computing $\mathcal{S}$ to compute the value of the appropriate iterate of the operation defining $\Phi$ applied to the value being carried by that wire in the circuit. We claim that this can be computed efficiently via dynamic programming. We start with the wires at the beginning of the circuit and work our way towards the end. It suffices to verify that if we know the appropriate values for the input wires, we can efficiently compute the corresponding value of the output wire. This involves the operations of tensor summation, multiplication of a tensor by a scalar sum, and tensor product of tensor with a vector in the last coordinate. It is straightforward to see that any of the allowed operations in $\mathcal{S}$ are compatible with this computation and that it runs in polynomial time in the relevant parameters. ∎
@@ 3.2 Algorithm and Analysis
1. Let $N$ be a positive integer quantifying the number of samples drawn from the relevant distributions in each step. 2. Let $A_{1}$ be the average of $N$ runs of $\mathcal{S}_{2}$ on $N$ independent samples drawn from $\mathcal{D}_{2}$, thought of as a $d\times d$ matrix. 3. Let $W_{1}$ be the span of the top-$k$ singular vectors of $A_{1}$. Let $\phi_{1}:\mathbb{R}^{1}\otimes\mathbb{R}^{d}\to\mathbb{R}^{k}$ be the projection of $\mathbb{R}^{d}$ onto $W_{1}$ composed with an isometry from $W_{1}$ to $\mathbb{R}^{k}$. 4. For $r=1$ to $m-1$: (a) Let $\Phi_{r}:(\mathbb{R}^{d})^{\otimes r}\to\mathbb{R}^{k}$ be the order-$r$ recursive pseudo-projection given by $\phi_{1},\ldots,\phi_{r}$. (b) Let $\Phi^{\prime}_{r}:(\mathbb{R}^{d})^{\otimes(2r+2)}\to\mathbb{R}^{k^{2}d^{2}}$ be the recursive pseudo-projection $\Phi_{r}\otimes I_{d}\otimes\Phi_{r}\otimes I_{d}$. (c) Let $A_{r+1}$ be the average of $N$ copies of $\Phi_{r}^{\prime}$ applied to the outputs of $\mathcal{S}_{2r+2}$ on $N$ independent samples drawn from $\mathcal{D}_{2r+2}$, using Lemma 3.4 to compute efficiently, thought of as a $dk\times dk$ matrix. (d) Let $W_{r+1}\subset\mathbb{R}^{dk}$ be the span of the top-$k$ singular vectors of $A_{r+1}$. (e) Let $\phi_{r+1}$ be the composition of the projection of $\mathbb{R}^{k}\otimes\mathbb{R}^{d}\to W_{r+1}$ with an isometry from $W_{r+1}$ to $\mathbb{R}^{k}$. 5. Let $\Phi_{m}:(\mathbb{R}^{d})^{\otimes m}\to\mathbb{R}^{k}$ be the recursive pseudo-projection given by $\phi_{1},\ldots,\phi_{m}$. 6. Let $A$ be the average of $N$ copies of $\Phi_{m}$ applied to the output of $\mathcal{S}_{m}$ on $N$ independent samples drawn from $\mathcal{D}_{m}$, using Lemma 3.4 to compute efficiently. 7. Let $B(\mathcal{X})$ be the average of $N$ copies of $\Phi_{m}$ applied to the output of $\mathcal{F}_{m}$ on the single sample $\mathcal{X}$ and $N$ independent samples from $\mathcal{D}^{\prime}$, using Lemma 3.4 to compute efficiently. Return $\langle A,B(\mathcal{X})\rangle$.
@@
We will establish the following lemma:
@@ 4.1 Technical Machinery
4.2 Learning Positive Linear Combinations of ReLU Activations
@@
Definition 4.5 (Positive Linear Combinations of ReLUs).
Let $\mathcal{C}_{d,k}$ denote the concept class of one-hidden-layer ReLU networks on $\mathbb{R}^{d}$ with $k$ hidden units and positive coefficients. That is, $F\in\mathcal{C}_{d,k}$ if and only if there exist $k$ unit vectors $v_{i}\in\mathbb{R}^{d}$ and non-negative coefficients $w_{i}\in\mathbb{R}_{+}$, $i\in[k]$, with $\sum_{i=1}^{k}w_{i}=1$ such that $F(x)=\sum_{i=1}^{k}w_{i}\mathrm{ReLU}(v_{i}\cdot x)$, where $\mathrm{ReLU}(t)=\max\{0,t\}$, $t\in\mathbb{R}$.
Following prior work on this problem [DKKZ20, GGJ+20, DK20, CKM21, CGKM22, CDG+23, DK24], we will assume that the feature vectors are normally distributed. The definition of the PAC learning problem in our setting is the following.
Definition 4.6 (PAC Learning $\mathcal{C}_{d,k}$).
The PAC learning problem for the class $\mathcal{C}_{d,k}$ is the following: The input is a multiset of i.i.d. labeled examples $(x,y)$, where $x\sim N(0,I)$ and $y=F(x)$, for an unknown $F\in\mathcal{C}_{d,k}$. The goal of the learner is to output a hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ that with high probability is close to $F$ in $L_{2}$-norm, i.e., satisfies $\|H-F\|_{2}\leq\epsilon\|F\|_{2}$. The hypothesis $H$ is allowed to lie in any efficiently representable hypothesis class $\mathcal{H}$.
@@
Theorem 4.7 (Learning Algorithm for $\mathcal{C}_{d,k}$).
@@
There is a PAC learning algorithm for $\mathcal{C}_{d,k}$ with respect to the standard Gaussian distribution on $\mathbb{R}^{d}$ with the following performance guarantee: Given $\epsilon>0$, and access to labeled examples from an unknown target $F\in\mathcal{C}_{d,k}$, the algorithm has sample and computational complexity $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$, and outputs a hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ that with high probability satisfies $\|H-F\|_{2}\leq\epsilon\|F\|_{2}$.
@@
It is easy to see that our results straightforwardly extend to the case that the labels have been corrupted by random zero-mean additive noise.
@@
As will become clear from the analysis of this subsection, our PAC learning algorithm is not proper. The hypothesis $H$ output by our algorithm is a succinct description of a low-degree polynomial (namely, of degree $\mathrm{poly}(1/\epsilon)$). Specifically, each Hermite coefficient will be expressed in compressed form, as per Proposition 1.7. This succinct description allows us to produce an efficient evaluation oracle for $H(x)$.
To apply Proposition 1.7 for this learning problem, we will need the following basic facts on the Hermite decomposition of the relevant functions.
Hermite Decomposition of Functions in $\mathcal{C}_{d,k}$
@@
Lemma 4.8 (Hermite Expansion of $\mathcal{C}_{d,k}$).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be any function of the form $F(x)=\sum_{i=1}^{k}w_{i}\mathrm{ReLU}(v_{i}\cdot x)$ for some $v_{i}\in\mathbb{R}^{d}$ with $\|v_{i}\|_{2}=1$ and $w_{i}\in\mathbb{R}_{+}$. Then $F$ has Hermite expansion $F(x)=\sum_{n=0}^{\infty}\langle T_{n},H_{n}(x)\rangle$, where $H_{n}$ is the normalized Hermite tensor (Definition 2.2) and $T_{n}\in(\mathbb{R}^{d})^{\otimes n}$ is defined by
$T_{n}\stackrel{{\scriptstyle{\mathrm{\footnotesize def}}}}{{=}}c_{n}\sum_{i=1}^{k}w_{i}v_{i}^{\otimes n}\;,$ (4)
where $c_{n}$ are explicit constants defined as follows: If $n>1$, then $c_{n}=0$ if $n$ is odd and
$c_{n}=(-1/4)^{(n-2)/4}\sqrt{\binom{n-2}{(n-2)/2}}/\sqrt{2\pi n(n-1)}=\Theta(n^{-5/4})$
if $n$ is even.
@@
By orthogonality of the Hermite tensors, we have that the tensor $T_{n}$ in the Hermite expansion of $F$, $F(x)=\sum_{n=0}^{\infty}\langle T_{n},H_{n}(x)\rangle$, is defined by $T_{n}=\mathbf{E}_{X\sim N(0,I)}[F(X)H_{n}(X)]$.
Lemma 3.1 of [DK24] shows that for all $n\in\mathbb{Z}_{+}$, we have that $\mathbf{E}_{G\sim N(0,1)}[\mathrm{ReLU}(G)h_{n}(G)]=c_{n}$ for the values of $c_{n}$ in the lemma statement. This in turn implies that for any unit vector $v\in\mathbb{R}^{d}$, it holds that $\mathrm{ReLU}(v\cdot x)=\sum_{n=0}^{\infty}c_{n}\langle H_{n}(x),v^{\otimes n}\rangle$ (Lemma 3.2 of [DK24]). Via orthogonality, we additionally get that $\mathbf{E}_{X\sim N(0,I)}[\mathrm{ReLU}(v\cdot X)H_{n}(X)]=c_{n}v^{\otimes n}$ (Corollary 3.3 of [DK24]).
Since $F(x)=\sum_{i=1}^{k}w_{i}\mathrm{ReLU}(v_{i}\cdot x)$, by linearity of expectation it follows that
$\mathbf{E}_{X\sim N(0,I)}[F(X)H_{n}(X)]=\sum_{i=1}^{k}w_{i}\mathbf{E}_{X\sim N(0,I)}[\mathrm{ReLU}(v_{i}\cdot X)H_{n}(X)]=\sum_{i=1}^{k}w_{i}c_{n}v_{i}^{\otimes n}=c_{n}\sum_{i=1}^{k}w_{i}v_{i}^{\otimes n}\;.$
@@
The key properties that enable our algorithm are that the tensor $H_{n}(x,y)$ of Definition 4.1 (i) can be computed by a sequential tensor computation of size $O(n)$, (ii) it satisfies $\mathbf{E}_{Y\sim N(0,I)}[H_{n}(x,Y)]=H_{n}(x)$, for any $x\in\mathbb{R}^{d}$, and (iii) it has bounded second moment (Lemma 4.3).

### v2

@@ 3.2 Recursive Pseudo-projections and their Properties
In order to describe the subspaces $W_{t}$, we will need a general tool that we call a recursive pseudo-projection, which we develop below. We start by defining the notion of a pseudo-projection.
@@
At any point in the circuit defining $\mathcal{S}$ if there is an order-$s$ tensor, we can compute the first $s$ steps of $\Phi$ applied to that value. The goal is to for every wire in the circuit computing $\mathcal{S}$ to compute the value of the appropriate iterate of the operation defining $\Phi$ applied to the value being carried by that wire in the circuit. We claim that this can be computed efficiently via dynamic programming (see Figure 1 for an illustration). We start with the wires at the beginning of the circuit and work our way towards the end. It suffices to verify that if we know the appropriate values for the input wires, we can efficiently compute the corresponding value of the output wire. This involves the operations of tensor summation, multiplication of a tensor by a scalar sum, and tensor product of tensor with a vector in the last coordinate. It is straightforward to see that any of the allowed operations in $\mathcal{S}$ are compatible with this computation and that it runs in polynomial time in the relevant parameters. ∎
Figure 1: The upper circuit represents the original Sequential Tensor Computation (STC). The markings next to the circuit elements represent the complexity of the elements being dealt with. The original inputs are vectors (lines). Then many of the middle components deal with $2$-tensors (squares), and finally the circuit elements on the right deal with $3$-tensors (cubes). In an STC of higher degree, one might need to use even higher degree tensors which quickly become computationally unmanageable. Fortunately, we are able to simulate this computation using the circuit below, keeping track of the projected values of each circuit element. This keeps the computational complexity low by reducing everything to vectors of dimension at most $\dim(\Phi)$ (represented by short lines).
@@ 3.3 Algorithm and Analysis
Input: Parameters $k,d,m\in\mathbb{Z}_{+}$; STCs $\mathcal{S}_{t},\mathcal{F}_{m}$, with size bound $S\in\mathbb{Z}_{+}$ and second moment bound $V\in\mathbb{R}_{+}$; $N$ i.i.d. samples from $\mathcal{D}_{t}$, $\mathcal{D}^{\prime}$, single sample $\mathcal{X}\sim X$; failure probability $\tau$. Output: A real number approximating $\langle M_{m},T(\mathcal{X})\rangle$, where $M_{m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}$ and $T(\mathcal{X})=\mathbf{E}_{\mathcal{Y}\sim\mathcal{D^{\prime}}}[\mathcal{F}_{m}(\mathcal{X},\mathcal{Y})]$, as defined in (1). 1. Let $N$ be a positive integer quantifying the number of samples drawn from the relevant distributions in each step. 2. Let $A_{1}$ be the average of $N$ runs of $\mathcal{S}_{2}$ on $N$ independent samples drawn from $\mathcal{D}_{2}$, thought of as a $d\times d$ matrix. 3. Let $W_{1}$ be the span of the top-$k$ singular vectors of $A_{1}$. Let $\phi_{1}:\mathbb{R}^{1}\otimes\mathbb{R}^{d}\to\mathbb{R}^{k}$ be the projection of $\mathbb{R}^{d}$ onto $W_{1}$ composed with an isometry from $W_{1}$ to $\mathbb{R}^{k}$. 4. For $r=1$ to $m-1$: (a) Let $\Phi_{r}:(\mathbb{R}^{d})^{\otimes r}\to\mathbb{R}^{k}$ be the order-$r$ recursive pseudo-projection given by $\phi_{1},\ldots,\phi_{r}$. (b) Let $\Phi^{\prime}_{r}:(\mathbb{R}^{d})^{\otimes(2r+2)}\to\mathbb{R}^{k^{2}d^{2}}$ be the recursive pseudo-projection $\Phi_{r}\otimes I_{d}\otimes\Phi_{r}\otimes I_{d}$. (c) Let $A_{r+1}$ be the average of $N$ copies of $\Phi_{r}^{\prime}$ applied to the outputs of $\mathcal{S}_{2r+2}$ on $N$ independent samples drawn from $\mathcal{D}_{2r+2}$, using Lemma 3.7 to compute efficiently, thought of as a $dk\times dk$ matrix. (d) Let $W_{r+1}\subset\mathbb{R}^{dk}$ be the span of the top-$k$ singular vectors of $A_{r+1}$. (e) Let $\phi_{r+1}$ be the composition of the projection of $\mathbb{R}^{k}\otimes\mathbb{R}^{d}\to W_{r+1}$ with an isometry from $W_{r+1}$ to $\mathbb{R}^{k}$. 5. Let $\Phi_{m}:(\mathbb{R}^{d})^{\otimes m}\to\mathbb{R}^{k}$ be the recursive pseudo-projection given by $\phi_{1},\ldots,\phi_{m}$. 6. Let $A$ be the average of $N$ copies of $\Phi_{m}$ applied to the output of $\mathcal{S}_{m}$ on $N$ independent samples drawn from $\mathcal{D}_{m}$, using Lemma 3.7 to compute efficiently. 7. Let $B(\mathcal{X})$ be the average of $N$ copies of $\Phi_{m}$ applied to the output of $\mathcal{F}_{m}$ on the single sample $\mathcal{X}$ and $N$ independent samples from $\mathcal{D}^{\prime}$, using Lemma 3.7 to compute efficiently. Return $\langle A,B(\mathcal{X})\rangle$.
@@
Controlling the size of the $\eta_{r}$ will be vital to our analysis. We begin by establishing the following recursive bound:
@@ 4.1 Technical Machinery
4.2 Learning Positive Linear Combinations of Non-Linear Activations
@@
Definition 4.5 (Positive Linear Combinations of an Activation Function).
Let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ be an activation function. Denote by $\mathcal{C}_{\sigma,d,k}$ the class of functions on $\mathbb{R}^{d}$ given by a one-layer network with positive coefficients using the activation $\sigma$. In particular, a function $F\in\mathcal{C}_{\sigma,d,k}$ if and only if there exist $k$ unit vectors $v_{i}\in\mathbb{R}^{d}$ and non-negative coefficients $w_{i}\in\mathbb{R}_{+}$, $i\in[k]$, with $\sum_{i=1}^{k}w_{i}=1$ such that $F(x)=\sum_{i=1}^{k}w_{i}\sigma(v_{i}\cdot x)$.
This is a prototypical family of neural networks whose learnability has been extensively studied over the past decade; see, e.g., [SJA16, ZSJ+17, GLM18, BJW19, DKKZ20, DK20, CKM21]. A particularly noteworthy special case is that of ReLU networks, where $\sigma(u)=\mathrm{ReLU}(u):=\max(0,u).$ As the majority of prior work on this problem, we will assume that the feature vectors are normally distributed. The definition of the PAC learning problem in our setting is the following.
Definition 4.6 (PAC Learning $\mathcal{C}_{\sigma,d,k}$).
The PAC learning problem for the class $\mathcal{C}_{\sigma,d,k}$ is the following: The input is a multiset of i.i.d. labeled examples $(x,y)$, where $x\sim N(0,I)$ and $y=F(x)$, for an unknown $F\in\mathcal{C}_{\sigma,d,k}$, for a known activation $\sigma$. The goal of the learner is to output a hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ that with high probability is close to $F$ in $L_{2}$-norm, i.e., satisfies $\|H-F\|_{2}\leq\epsilon$. The hypothesis $H$ is allowed to lie in any efficiently representable hypothesis class $\mathcal{H}$.
@@
Theorem 4.7 (Learning Algorithm for $\mathcal{C}_{\sigma,d,k}$).
Suppose that the activation $\sigma$ has $\|\sigma\|_{4}=O(1)$. Letting $c_{\sigma,t}:=\mathbf{E}[\sigma(G)h_{t}(G)]$ be the $t^{th}$ Hermite coefficient of $\sigma$, suppose additionally that for some $\epsilon,\delta>0$ and some positive $n$ it holds:
  • $\sum_{t>n}c_{\sigma,t}^{2}<\epsilon^{2}/4$
  • $|c_{\sigma,t}|\geq\delta$ for all $1\leq t\leq 2n$ unless $t$ is odd and $c_{\sigma,t}=0$.
@@
Then there is a PAC learning algorithm for $\mathcal{C}_{\sigma,d,k}$ with respect to the standard Gaussian distribution on $\mathbb{R}^{d}$ with the following performance guarantee: Given $\epsilon>0$, $\sigma,k,d$ and access to labeled examples from an unknown target $F\in\mathcal{C}_{\sigma,d,k}$, the algorithm has sample and computational complexity $\mathrm{poly}(dk/(\epsilon\delta))2^{O(n)}$, and outputs an efficiently computable hypothesis $H:\mathbb{R}^{d}\to\mathbb{R}$ that with high probability satisfies $\|H-F\|_{2}\leq\epsilon$.
@@
It is easy to see that the algorithm of Theorem 4.7 straightforwardly extends to the case that the labels have been corrupted by random zero-mean additive noise.
@@
As will become clear from the analysis that follows, our PAC learning algorithm is not proper. The hypothesis $H$ it outputs is a succinct description of a low-degree polynomial (namely, of degree $n$). Specifically, each Hermite coefficient will be expressed in compressed form, as per Proposition 3.2. This succinct description allows us to produce an efficient evaluation oracle for $H(x)$.
Remark 4.8.
Theorem 4.7 should provide useful results for most activation functions $\sigma$ that are not odd (which would lead to vanishing of the even degree coefficients). Unfortunately, for many such activations, the analysis will be somewhat subtle—as one needs to show that $c_{\sigma,t}$ is not too close to 00, which usually can only be done by computing these Hermite coefficients exactly. The two examples given below correspond to common activations for which this could be easily done.
In particular, for the special case of sums of ReLUs we obtain the following:
Corollary 4.9.
For $\sigma(u)=\mathrm{ReLU}(u)$, there is a PAC learning algorithm that learns $\mathcal{C}_{\sigma,d,k}$ to error $\epsilon$ with sample and computational complexity $\mathrm{poly}(d,k)2^{\mathrm{poly}(1/\epsilon)}$.
Proof.
Given Theorem 4.7, this amounts to proving some basic statements about the Fourier coefficients of $\sigma$. In particular, it follows immediately from the claim that $c_{\sigma,t}=0$ if $t>1$ is odd and
$c_{\sigma,t}=(-1/4)^{(t-2)/4}\sqrt{\binom{t-2}{(t-2)/2}}/\sqrt{2\pi t(t-1)}=\Theta(t^{-5/4})$
if $t$ is even. This follows from Lemma 3.1 of [DK24]. Thus, we may take $n=O(\epsilon^{-2/3})$ and the rest follows. ∎
Another application is to periodic activations given by $\sigma(u)=\cos(\gamma u)$ for some frequency parameter $\gamma$. While the majority of work on neural networks considers monotonic activations (e.g., ReLUs), an important branch of the literature considers networks with periodic activations, specifically the cosine activation. Such activations are particularly useful in signal processing and computer vision applications, where it has been empirically observed that networks with periodic activations are capable of representing details in the signals better than ReLU networks. The reader is referred to  [SMB+20, MST+20, RNM+21, VPHA24] and references therein.
In particular, we show:
Corollary 4.10.
For $\sigma(u)=\cos(\gamma u)$, for some parameter $\gamma>0$, there is a PAC learning algorithm that learns $\mathcal{C}_{\sigma,d,k}$ to error $\epsilon$ with sample and computational complexity $2^{O(1/\gamma^{2})}\mathrm{poly}(dk/\epsilon)$.
While the complexity of our algorithm blows up as $\gamma$ goes to infinity, there is evidence that this kind of dependence is necessary even for $k=1$. In particular, [DKRS23] (see Theorem 1.13) implies that any Statistical Query (SQ) algorithm learning a single cosine activation to non-trivial error requires $\exp(\min(\Omega(\gamma^{2}),d^{\Omega(1)}))$ resources. Furthermore,  [SZB21] shows that if a tiny amount of adversarial label noise is added, then the problem is hard for polynomially-sized $\gamma$ under certain cryptographic assumptions. Note that for $\gamma\ll 1/\sqrt{\log(dk/\epsilon)}$, our algorithm runs in $\mathrm{poly}(dk/\epsilon)$ time. By the above discussion, this is best possible for SQ algorithms.
Proof of Corollary 4.10.
In light of Theorem 4.7, this result comes down to calculating $c_{\sigma,t}$. Note that $c_{\sigma,t}$ is the real part of
[…]
In order to analyze this, we begin by computing
[…]
for real valued $\alpha$. By completing the square, we see that this is
[…]
Note that by Lemma 2.7 of [Kan21] (and noting the difference in normalization between their Hermite polynomials and ours), we have that
$\mathbf{E}[h_{t}(G+\alpha)]=\alpha^{t}/\sqrt{t!}.$
Thus, we have that for real $\alpha$
[…]
However, by analytic continuation, this must also hold for all complex $\alpha$ as well. Thus, plugging in $\alpha=i\gamma$ and taking the real part we have that
$c_{\sigma,t}=\begin{cases}e^{-\gamma^{2}/2}\gamma^{t}(-1)^{t/2}/\sqrt{t!}&\textrm{if }t\textrm{ is even}\\ 0&\textrm{if }t\textrm{ is odd}.\end{cases}$
From this, it is not hard to see that it suffices to take $n=O(1/\gamma^{2}+\log(1/\epsilon))$ and $\delta=\Omega(e^{-\gamma^{2}}\epsilon^{2})$ in Theorem 4.7, yielding our result. ∎
Remark 4.11.
Corollary 4.10 easily generalizes to periodic activations $\sigma(u)=\cos(\gamma u+\theta)$ with an additional phase $\theta$, so long as $\theta$ is not close to a multiple of $\pi/2$ (which would again lead to vanishing Fourier coefficients).
To prove Theorem 4.7, our basic strategy is to study the Hermite expansion of $F$, which can be related to the kind of tensors approximated in Proposition 3.2. Our hypothesis will then be an approximation to the low degree part of the Hermite expansion of $F$. To make this work, we will first need to understand the Hermite expansion of the relevant functions.
Hermite Decomposition of Functions in $\mathcal{C}_{\sigma,d,k}$
@@
Lemma 4.12 (Hermite Expansion of $\mathcal{C}_{\sigma,d,k}$).
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be any function of the form $F(x)=\sum_{i=1}^{k}w_{i}\sigma(v_{i}\cdot x)$ for some $v_{i}\in\mathbb{R}^{d}$ with $\|v_{i}\|_{2}=1$ and $w_{i}\in\mathbb{R}_{+}$. Then $F$ has Hermite expansion $F(x)=\sum_{m=0}^{\infty}\langle T_{m},H_{m}(x)\rangle$, where $H_{m}$ is the normalized Hermite tensor (Definition 2.2) and $T_{m}\in(\mathbb{R}^{d})^{\otimes n}$ is defined by
$T_{m}\stackrel{{\scriptstyle{\mathrm{\footnotesize def}}}}{{=}}c_{\sigma,m}\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}.$ (6)
@@
By orthogonality of the Hermite tensors, we have that the tensor $T_{n}$ in the Hermite expansion of $F$, $F(x)=\sum_{m=0}^{\infty}\langle T_{m},H_{m}(x)\rangle$, is defined by $T_{m}=\mathbf{E}_{X\sim N(0,I)}[F(X)H_{m}(X)]$.
By definition, $\mathbf{E}_{G\sim N(0,1)}[\sigma(G)h_{m}(G)]=c_{\sigma,m}$. This in turn implies that for any unit vector $v\in\mathbb{R}^{d}$, it holds that $\sigma(v\cdot x)=\sum_{m=0}^{\infty}c_{\sigma,m}\langle H_{m}(x),v^{\otimes m}\rangle$ (Lemma 3.2 of [DK24]). Via orthogonality, we additionally get that $\mathbf{E}_{X\sim N(0,I)}[\sigma(v\cdot X)H_{m}(X)]=c_{\sigma,m}v^{\otimes m}$ (Corollary 3.3 of [DK24]).
Since $F(x)=\sum_{i=1}^{k}w_{i}\sigma(v_{i}\cdot x)$, by linearity of expectation it follows that
$\mathbf{E}_{X\sim N(0,I)}[F(X)H_{m}(X)]=\sum_{i=1}^{k}w_{i}\mathbf{E}_{X\sim N(0,I)}[\sigma(v_{i}\cdot X)H_{m}(X)]=\sum_{i=1}^{k}w_{i}c_{\sigma,m}v_{i}^{\otimes m}=c_{\sigma,m}\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}\;.$
@@
The key properties that enable our algorithm are that the tensor $H_{m}(x,y)$ of Definition 4.1 (i) can be computed by a sequential tensor computation of size $O(m)$, (ii) it satisfies $\mathbf{E}_{Y\sim N(0,I)}[H_{m}(x,Y)]=H_{m}(x)$, for any $x\in\mathbb{R}^{d}$, and (iii) it has bounded second moment (Lemma 4.3).
@@
To apply Proposition 3.2, we need to relate $T_{m}$ (which we can approximate) to the standard moment tensors, $M_{m}$. Fortunately, it follows from the definitions that $M_{t}=T_{t}/c_{\sigma,t}$. By assumption, for $t\leq 2n$ even, we have $|c_{\sigma,t}|>\delta$ and so we can approximate $M_{t}$ to error $\eta$ by approximating $T_{t}$ to error $\eta/\delta$. If $t$ is odd and $c_{\sigma,t}=0$, we will have no need to estimate $M_{t}$. Otherwise, for $t$ odd we can approximate $M_{m}$ to error $\eta$ by approximating $T_{t}$ to error $\eta/\delta$ again.